# Leveraging Domain Knowledge for Efficient Reward Modelling in RLHF: A Case-Study in E-Commerce Opinion Summarization

### v1

@@ 3.3 OpinPref Dataset
We create OpinPref by asking humans to rate opinion summaries for given reviews. We utilize domain experts (annotator details in Appendix G) to perform the annotation. We observe a Fleiss’ Kappa ($\kappa$) score of $62.67\%$ (substantial aggrement; aggrement is substantial when $60\%\leq\kappa<80\%$). We believe that aligning to the internal reward model of domain experts would lead to better opinion summaries. We provide the domain expert with reviews of a product and two opinion summaries (products are sampled from the PromptOpinSumm dataset). The domain expert notifies which of the two summaries they prefer. We use this to construct a dataset of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We construct a dataset of $940$ samples. Appendix F includes statistics on the dataset (such as length of reviews and summaries).
@@
Figure 2: Overview of our approach. Step-$1$: We generate a new dataset for training Opinion Summarizers: PromptOpinSumm, by prompting Mistral-$7$B  model. Again, we use Mistral-$7$B  to compute values for the $7$ features discussed in Section 4.1. Step-$2$: We ask humans (domain experts) to given their preference, given reviews and two opinion summaries (A, B). We use the preference data to train the reward model, $\varphi_{op}$. Step-$3$: We sample instances from PromptOpinSumm dataset; $\varphi_{op}$ assigns a score to the sampled summaries, the policy, $\pi_{\theta}$, assigns log probabilities to these summaries. Proximal Policy Optimization uses these to update $\pi_{\theta}$.
4 Technique
@@
We test our hypothesis in the domain of Opinion Summarization in E-Commerce – the task of summarizing user reviews for a product. Typically, user reviews discuss several aspects of a product, opinions/sentiments towards these aspects. An opinion summary has to reflect all the aspects that the input reviews discuss, along with the opinions/sentiments expressed towards these aspects. The section is structured as follows: in Section 4.1, we discuss our reward modelling technique, human annotation, etc., and in Section 4.2, we discuss the RLHF training pipeline using the novel reward model.
4.1 Domain Knowledge Induced Reward Modelling
@@
We leverage insights on desirable properties in an opinion summary from domain experts [^5] We consulted domain experts from an E-Commerce platform.. Based on these insights, we characterize our reward model, $\varphi_{op}$, as follows: $\varphi_{op}$$\>=f(v)$, where $v\in$aspect-coverage, opinion-faithfulness, opinion-coverage, conciseness, relevance, hallucination, language-correctness}. The features aspect-coverage, opinion-faithfulness  and opinion-coverage  are directly relevant to Opinion Summarization. They check if the generated opinion summary covers all mentioned aspects and opinions. The features conciseness, relevance  and hallucination  pertain to Generic Text Summarization Rush et al. (2015). They check if the generated summary is concise, relevant to the input reviews, and is free from hallucination. The features language-correctness  pertains to the grammar of the language, checking if the generated text follows the rules of the language. These features, together, characterize the goodness of an opinion summary. We use Mistral-$7$B (see Appendix A for details) to obtain values for these features for an opinion summary, given reviews. We denote this transformation using $\Phi$.
@@
We train $\varphi_{op}$ using OpinPref, which is of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We parameterize $\varphi_{op}$ using a Feed-Forward Neural Network, and train it using the Elo-loss Ouyang et al. (2022); Glaese et al. (2022) (Equation 4.1; $\Phi(R,s_{i})$ uses Mistral-$7$B to compute the $7$ features). Such a formulation for $\varphi_{op}$ brings interpretability – which features influence human preference the most, and is free from Alignment Tax Bai et al. (2022a) – degradation of LM on benchmarks after reward model training.
@@
4.2 RLHF Training Pipeline
Using the trained reward model, we follow a similar training pipeline as Bai et al. (2022a); Ouyang et al. (2022), with a modification: Limited Trajectory Reinforcement Learning. Computing the transformation $\Phi$ for each generation online (during training) is expensive, especially with limited compute resources. To circumvent this, we limit the trajectories that are explored by our policy, $\pi_{\theta}$. Specifically, we limit it to the Good-Sum, SBad-Sum and VBad-Sum trajectories in the PromptOpinSumm dataset. Having varying levels of quality in PromptOpinSumm is of use here – it lets the model still explore trajectories of several quality. Thus, we have an offline experience buffer, with $\Phi$ precomputed, for $\pi_{\theta}$ learn from.
We use Proximal Policy Optimization (PPO) Schulman et al. (2017) to train our model (Equation 4.2). For each training step, we sample ($R,s,\Phi(R,s)$) tuples from PromptOpinSumm. We use the trained $\varphi_{op}$ to compute the reward for $s$ ($=\varphi_{op}\big{(}\Phi(R,s)\big{)}$). PPO uses this to update the log probability assigned by $\pi_{\theta}$. We parameterize $\pi_{\theta}$ using a Transformer model, which takes reviews as input, and generates an opinion summary.
Model-Code Amazon Flipkart Oposum+
@@ 4.2 RLHF Training Pipeline
Prior Works MeanSum Chu and Liu (2018) $29.20$ $4.70$ $18.15$- -- -- $26.25$ $4.62$ $16.49$
CopyCat Bražinskas et al. (2020) $31.97$ $5.81$ $20.16$- -- -- $27.98$ $5.79$ $17.07$
PlanSum Amplayo and Lapata (2020) $32.87$ $6.12$ $19.05$- -- -- $30.26$ $5.29$ $17.48$
MultimodalSum Im et al. (2021) $34.19$ $7.05$ $20.81$- -- -- $33.08$ $7.46$ $19.75$
Op-Sum-GenSiledar et al. (2023a) $35.46$ $7.30$ $21.50$- -- -- $36.57$ $8.79$ $21.35$
Ours’ Supervised $28.22$ $4.91$ $16.79$ $28.10$ $4.18$ $15.37$ $30.15$ $7.26$ $16.74$
NaiveMean $28.09$ $4.89$ $16.75$ $26.42$ $4.14$ $15.0$ $30.62$ $7.62$ $17.11$
Synth-Feedback $24.82$ $4.30$ $15.48$ $27.40$ $3.64$ $15.25$ $29.10$ $6.19$ $16.20$
Inductive-Bias $29.16$ $4.74$ $17.07$ $27.14$ $4.0$ $15.0$ $30.75$ $7.45$ $16.82$
Table 3: Overlap-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$). Our models are not near the state-of-the-art in these metrics. We expect such a trend, and we present the reasons in Section 3.2. We include these results for completeness. Note that for Flipkart benchmark, we do not have results from Op-Sum-Gen, as Siledar et al. (2023a) only provide aspect specific summarization model for the Flipkart benchmark.
5 Experiments and Results
Empirically, we test our technique exhaustively against the current State-of-the-Art model, and strong RL and RLHF baselines (our own design and from contemporary literature). We find that in overlap-based evaluation (Rouge-$1$, Rouge-$2$, Rouge-$L$), our proposed model (and other RL baselines) falls short of the SOTA. However, this is expected due to several shortcomings (Section 3.2) in the benchmark datasets. We conduct human and Gpt-$4$ evaluations to reliably verify efficacy of our technique. We find that our proposed technique excels significantly. In the rest of the section, we describe our models (Section 5.1) and evaluation results (Section 5.2).
5.1 Models
We train the following models on PromptOpinSumm:
NaiveMean: This is a Reinforcement Learning model, where the reward is computed by averaging the metric scores obtained using $\Phi$.
[…]
Inductive-Bias: This is a RLHF model, trained following our hypothesis. We train $\varphi_{op}$ using OpinPref dataset.
In addition to the models above, we also use the current SOTA Siledar et al. (2023a) for human evaluations. Siledar et al. (2023a) propose opinion summarizer models for general opinion summarization and aspect specific opinion summarization. As all of our models are for general opinion summarization, we use the same from Siledar et al. (2023a). We refer to this model by the name Op-Sum-Gen in the rest of the paper. Note that we do not use Direct Preference Optimization (DPO) Rafailov et al. (2023) and vanilla RLHF Ziegler et al. (2019); Bai et al. (2022a) as baselines. This is because both these techniques require huge human preference data. Additionally, the goal of the paper is not to propose a new RLHF technique, rather to propose a way to achieve alignment with domain-specific human goals with modest human preference annotation, using RLHF.
We use BART-Large Lewis et al. (2020) for all of our models. The choice of the model is governed by two factors: (a) It provides a similar environment (model size) for comparison with SOTA, (b) We find that LLMs (Mistral-$7$B, LLaMA2222-$7$B, Zephyr-$7$B, etc.) are already quite good at opinion summarization, thus any performance benefits (over SOTA) in those models cannot be reliably attributed to our novel reward modelling technique. We perform a hyperparameter sweep on batch size, learning rate, and learning rate warmup while training our models. We choose the model which performs best on the validation set of PromptOpinSumm to generate summaries for the benchmark datasets. We include more details on implementation in Appendix C.
@@ 5.2 Evaluation Results
On the models above, we perform overlap-based evaluations (Rouge). We report the scores in Table 3. In overlap-based evaluations, we find that our models do not perform better than existing models. We expect this behaviour (Section 3.2), which is why we do not draw conclusions from this evaluation. We report the metrics for completeness.
We also perform Gpt-$4$ and human evaluations. For human evaluations, we use domain experts (annotator details in Appendix G) for the evaluations; we observe a Fleiss’ Kappa ($\kappa$) score of $56.25\%$ (moderate agreement; agreement is moderate when $40\%\leq\kappa<60\%$). In both these settings, we ask the evaluator (Gpt-$4$ or domain expert) to rank the generated summaries given the reviews. We anonymize all the summaries and shuffle the orders for each evaluation instance. Given the rankings, we compute the fraction of pairwise wins, ties and losses among all the models. We compare summaries from Supervised, NaiveMean, Synth-Feedback, Inductive-Bias, Op-Sum-Gen models and ground truth summaries. We include ground truth summaries in the evaluation to verify our claims about the quality of the benchmarks. Figure 1 shows the performance of Inductive-Bias model over the others for human evaluations for the Amazon benchmark. We see that our proposed technique (domain knowledge based reward model $\varphi_{op}$) helps Inductive-Bias model achieve summaries which are always preferred more than the competitors. We can see a similar trend for Gpt-$4$ evaluation too, for all the benchmarks (Figures 3, 4 and 5). Both these evaluations empirically show that RLHF, using modest preference annotation, through domain knowledge infusion, can lead to performance gains.
@@
Figure 4: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for the Flipkart benchmark. Note that for Flipkart benchmark, we do not have results from Op-Sum-Gen, as Siledar et al. (2023a) only provide aspect specific summarization model for Flipkart benchmark.
@@
(a) aspect-coverage
(b) opinion-faithfulness
(c) relevance
(d) hallucination
Figure 6: Scores on domain knowledge based features on the Amazon benchmark for top-$3$ models (IB: Inductive-Bias, NM: NaiveMean, SF: Synth-Feedback). We see that Inductive-Bias model manages to stay ahead of the competitors, notably for hallucination. This increases our confidence on the Inductive-Bias model, as summaries generated by this model align more with the domain features.

### v2

@@ 3.3 OpinPref Dataset
We create OpinPref by asking humans to rank opinion summaries for given reviews. We utilize domain experts (annotator details in Appendix I) to perform the annotation. We believe that aligning with the internal reward model of domain experts would lead to better opinion summaries. We provide the domain expert with product reviews and two opinion summaries (products are sampled from the PromptOpinSumm dataset). The domain expert notifies which of the two summaries they prefer. We use this to construct a dataset of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We construct a dataset of $940$ samples. We observe a Fleiss’ Kappa ($\kappa$) score of $62.67\%$ (substantial aggrement; aggrement is substantial when $60\%\leq\kappa<80\%$). Appendix H includes statistics on the dataset.
@@
Figure 2: Overview of our approach. Step-$1$: We generate a new dataset for training Opinion Summarizers: PromptOpinSumm, by prompting Mistral-$7$B  model. Again, we use Mistral-$7$B  to compute values for the $7$ features discussed in Section 4.1. Step-$2$: We ask humans (domain experts) for their preference, given reviews and two opinion summaries (A, B). We use the preference data and the features to train the reward model, $\varphi_{op}$. Step-$3$: We sample instances from PromptOpinSumm dataset; $\varphi_{op}$ assigns a score to the sampled summaries, the policy, $\pi_{\theta}$, assigns log probabilities to these summaries. Proximal Policy Optimization uses these to update $\pi_{\theta}$.
4 Efficient Reward Modeling
We highlighted in Section 1 how the reward model ($\varphi$) can depend on the downstream task. Such dependence necessitates task/domain-specific human preference datasets, which are costly and time-consuming to create. This creates an obstacle in employing RLHF in task/domain-specific setups, thus hindering the steering of LLMs towards task/domain-specific human values.
We solve this challenge by leveraging domain knowledge. The key insight is that we can use the domain knowledge to impart some inductive biases into the mathematical modeling of $\mathbf{\varphi}$. This would significantly reduce the amount of data required for training $\varphi$. Specifically, we say that $\varphi_{\tau}$ (reward model for a domain $\tau$) can be modelled by some numeric features $v_{1},v_{2},\cdots,v_{n}$. These $n$ features fully characterize [^7] Example of such characterization: Features like fluency, coherence, etc. can characterize text generated by an LLM. the outputs from the LLM on some input. Thus, instead of training $\varphi_{\tau}$ on the text data ($\{(X,Y_{w},Y_{l})\>|\>Y_{w}\succ Y_{l}\}$), we use the $n$ features. Such a formulation for $\varphi$ also brings interpretability and frees $\varphi$ from Alignment Tax.
@@ 4 Efficient Reward Modeling
In Section 4.1, we detail our technique for the task/domain of E-Commerce Opinion Summarization—the task of summarizing user reviews for a product. Typically, user reviews discuss several aspects of a product and opinions/sentiments towards these aspects. An opinion summary must reflect all the aspects discussed by the input reviews and the opinions expressed towards these aspects. We discuss how we leverage such desirable properties to model $\varphi$.
4.1 Inducing Domain Knowledge
@@
We identify desirable properties in an opinion summary with the help of domain experts [^8] Domain experts are from an E-Commerce platform.. We held multiple discussions to finalize the set of desirable properties. We show that these properties are correlated to humans’ judgement of summary in Appendix A (Table 3). Based on these properties, we model $\varphi_{op}$ (reward model for opinion summarization) as: $\varphi_{op}$$\>=f(v)$, where $v\in$aspect-coverage, opinion-faithfulness, opinion-coverage, conciseness, relevance, hallucination, language-correctness}. The features aspect-coverage, opinion-faithfulness  and opinion-coverage  check if the generated opinion summary covers all mentioned aspects and opinions faithfully. The features conciseness, relevance , and hallucination  check if the generated summary is concise, relevant to the input reviews, and is free from hallucination. The feature language-correctness  checks if the generated text follows the language rules. We provide more details in Appendix A. These features, together, characterize the goodness of an opinion summary. We instruct Mistral-$7$B (Appendix A) to generate values for these features for an opinion summary, given reviews. We denote this transformation (from reviews and summary to $7$ features) using $\Phi$.
@@
We train $\varphi_{op}$ using OpinPref, which is of the form: $\mathcal{D}_{h}=\{(R,s_{w},s_{l})\>|\>s_{w}\succ s_{l}\}$, where $R$ is the set of reviews and $s_{w}$ and $s_{l}$ are opinion summaries. We parameterize $\varphi_{op}$ using a Feed-Forward Neural Network and train it using the Elo-loss Ouyang et al. (2022); Glaese et al. (2022) (Equation 4.1; $\Phi(R,s_{i})$ uses Mistral-$7$B to compute the $7$ features; only $\varphi_{op}$ is trainable, $\Phi$ is not).
After such an efficient reward modeling, we use $\varphi_{op}$ for regular RLHF training (Appendix B) to get an Opinion Summarizer aligned with human goals. We illustrate the whole flow in Figure 2.
@@
5 Experiments
We test our technique against the State-of-the-Art (SOTA) models, and strong Reinforcement Learning (RL) and RLHF baselines (our design and contemporary works). We list the questions we attempt to answer (through the experiments) in Section 5.1. We conduct automatic, human, and Gpt-$4$ evaluations to verify our claim. We find that our proposed technique excels significantly. In the rest of the section, we describe our models (Section 5.1) and evaluation results (Section 5.2).
5.1 Models & Objectives
We train the following models:
NaiveMean: This is a Reinforcement Learning model, where the reward is computed by averaging the feature values obtained using $\Phi$.
[…]
Inductive-Bias: This RLHF model is trained following our hypothesis (infusing domain knowledge into $\varphi$ to reap benefits of RLHF with modest human preference data). We train $\varphi_{op}$ using OpinPref dataset.
With these models, we ask the following questions in our experiments:
Scene-I: How effective is our technique (infusing domain knowledge into $\varphi$ to reap benefits of RLHF with modest human preference data) over and above the usage of a good training dataset? A comparative evaluation of Supervised and Inductive-Bias would answer this.
Scene-II: How effective is our technique over and above vanilla RL? A comparative evaluation of NaiveMean and Inductive-Bias would answer this.
Scene-III: How effective is our technique over contemporary RLHF techniques, which work without preference data? A comparative evaluation of Synth-Feedback and Inductive-Bias would answer this.
Scene-IV: How effective is our technique, agnostic of the preference data? This question is raised to answer whether the gains are solely due to the good quality of OpinPref, or the approach. A comparative evaluation between DPO (Rafailov et al. (2023), which uses OpinPref in a supervised fashion) and Inductive-Bias would answer this.
In addition to the above questions, we also check how our models fare against the SOTA (Op-Sum-Gen: Siledar et al. (2023a), Medos: Siledar et al. (2024), etc.). We do not use vanilla RLHF Ziegler et al. (2019); Bai et al. (2022a) as a baseline, as it requires huge human preference data. Given that the goal of the paper is not to propose a new RLHF technique, but rather to propose a way to use RLHF with modest human preference annotations, omitting vanilla RLHF as a baseline does not affect our conclusions in any way.
We use BART-Large Lewis et al. (2020) for all of our models. The choice of the model is governed by two factors: (a) It provides a similar environment (model size) for comparison with SOTA, (b) We find that LLMs (Mistral-$7$B, LLaMA2222-$7$B, Zephyr-$7$B, etc.) are already quite good at opinion summarization; thus any performance benefits (over SOTA) in those models cannot be reliably attributed to our approach. We include implementation details in Appendix E.
Model-Code Amazon Amazon-R Amazon-Rdq
@@ 5.1 Models & Objectives
Prior Works MeanSum Chu and Liu (2018) $29.20$ $4.70$ $18.15$- -- -- -- -- --
CopyCat Bražinskas et al. (2020) $31.97$ $5.81$ $20.16$ $20.09$ $1.79$ $12.94$ $20.54$ $1.94$ $13.85$
PlanSum Amplayo and Lapata (2020) $32.87$ $6.12$ $19.05$ $20.49$ $1.76$ $12.44$ $19.09$ $1.58$ $12.02$
MultimodalSum Im et al. (2021) $34.19$ $7.05$ $20.81$ $21.43$ $1.58$ $13.20$ $20.39$ $2.08$ $12.83$
Op-Sum-GenSiledar et al. (2023a) $\mathbf{35.46}$ $\underline{7.30}$ $\mathbf{21.50}$- -- -- -- -- --
MedosSiledar et al. (2024) $\underline{34.63}$ $\mathbf{7.48}$ $\underline{20.97}$ $23.92$ $2.27$ $14.69$ $25.44$ $4.16$ $16.45$
Ours’ DPO $23.96$ $4.54$ $14.27$ $26.37$ $4.25$ $15.03$ $25.13$ $3.84$ $14.86$
Supervised $28.99$ $4.90$ $16.91$ $32.52$ $5.96$ $18.07$ $30.46$ $\underline{5.49}$ $17.63$
NaiveMean $28.08$ $4.81$ $16.77$ $\mathbf{34.0}$ $\underline{6.30}$ $\underline{18.81}$ $\underline{30.97}$ $5.25$ $\underline{18.36}$
Synth-Feedback $29.39$ $4.68$ $17.35$ $33.62$ $6.06$ $18.61$ $30.65$ $5.23$ $18.11$
Inductive-Bias $28.41$ $4.65$ $16.90$ $\underline{33.95}$ $\mathbf{6.40}$ $\mathbf{19.23}$ $\mathbf{31.89}$ $\mathbf{5.78}$ $\mathbf{18.84}$
Table 1: Reference-based Evaluation Results (R-$1$: Rouge-$1$, R-$2$: Rouge-$2$, R-L: Rouge-$L$) for the Amazon, Amazon-R and Amazon-Rdq benchmarks. We see the following things: (a) Our proposed dataset (PromptOpinSumm) leads to marked increased over the SOTA (by $\sim 4$ R-L points), (b) Inductive-Biasproves to be the winner in all the four scenarios: Scene-I, Scene-II, Scene-III and Scene-IV (Section 5.1), proving the efficacy of our technique. We also see that for the Amazon benchmark, our models lag behind. However, this is expected, as we highlight in Section 3.2.
@@ 5.2 Evaluation Results
We test our approach on $\mathbf{9}$ benchmarks (Section 3.2). In the main manuscript, we report automatic evaluation results on Amazon-based benchmarks (Table 1), human evaluation on the Amazon benchmark, and Gpt-$4$ evaluations on Amazon, Flipkart and Oposum+ benchmarks. Liu et al. (2023) show that GPT-$4$ evaluations correlate well with human evaluations for summarization; hence, in the interest of time and monetary expense, we resort to Gpt-$4$ evaluation. We include automatic evaluation results for the rest of the benchmarks (Tables 4 and 5), and BertScore based evaluations (Table 6) for all the benchmarks in Appendix C. We also include model generations for a randomly sampled product in Appendix C (Table 14). Due to the shortcomings highlighted in Section 3.2 and Appendix D, we complement our automatic evaluations of Amazon, Flipkart and Oposum+ with human and Gpt-$4$ evaluations.
Automatic Evaluation.  From Table 1, we see that our proposed models are always better than the SOTA for Amazon-R and Amazon-Rdq. Supervised Fine Tuning on PromptOpinSumm (Supervised model) helps achieve significantly better Rouge scores. This highlights the efficacy of our proposed PromptOpinSumm dataset. From the automatic evaluations on Amazon-R and Amazon-Rdq, we see the following things:
Answer to Scene-I: We see that Inductive-Bias achieves gains over Supervised. This answers the question in Scene-I: Our technique is effective over and above using a good dataset.
Answer to Scene-II: We see that Inductive-Bias achieves gains over NaiveMean. This answers the question in Scene-II: Our technique is effective over vanilla RL.
Answer to Scene-III: We see that Inductive-Bias achieves gains over Synth-Feedback. This answers the question in Scene-III: Our technique is effective over the SOTA RLHF technique, which works without human preference data.
Answer to Scene-IV: We see that Inductive-Bias achieves gains over DPO. This verifies that gains of Inductive-Bias can be safely attributed to the approach (not just the quality of OpinPref).
Human/Gpt-$4$ Evaluation.  We conduct human evaluation (Figure 1) for the Amazon benchmark, using $3$ domain experts (details in Appendix I). We observe a Fleiss’ Kappa ($\kappa$) score of $56.25\%$ (moderate agreement; agreement is moderate when $40\%\leq\kappa<60\%$). We ask the experts to rank the summaries (anonymized and shuffled) given the reviews. Given the rankings, we compute the fraction of pairwise wins, ties, and losses among all the models. We compare summaries from Supervised, NaiveMean, Synth-Feedback, Inductive-Bias, Op-Sum-Gen (SOTA) models and ground truth summaries. We include ground truth summaries in the evaluation to verify our claims about the quality of the benchmarks. From Figure 1, we see that Inductive-Bias wins significantly over the competitors, further proving the efficacy of our technique.
We run Gpt-$4$ evaluations for Amazon, Flipkart and Oposum+ benchmarks (Figures 3, 4, 5). We run Gpt-$4$ evaluations for Amazon, as the agreement in human evaluation was moderate. We arrive at the same conclusions as human evaluation. We prompt Gpt-$4$ to rank the summaries (anonymized and shuffled) given the reviews. As before, we compute the fraction of wins, ties, and losses. Again, we see that Inductive-Bias remains a clear winner.
@@
Figure 4: Gpt-$4$ Eval: Pairwise win-tie-loss percentage of Inductive-Bias model vs. competitors, for Flipkart benchmark. Note that for the Flipkart benchmark, we do not have results from Op-Sum-Gen, as Siledar et al. (2023a) only provide aspect-specific summaries.